A car moves 46 seconds in 2 miles. Find the average speed.
Found 2 solutions by RAY100, stanbon:
Answer by RAY100(1637)

Velocity = Distance / Time
.
V= 2 miles / {46 sec * (1 hr/3600 sec ) } = 156.52 mi per hr

Answer by stanbon(75887)

A car moves 46 seconds in 2 miles. Find the average speed.
---
2/46 miles per second
---
(2/46 m/s)(3600 sec/hr) = (2*3600/46) mph = 156.52 mph
========================================================
Cheers,
Stan H.